OpenStudy (anonymous):

Please help me: Convert each of the following from decimal form to the given base. A) 65 to base seven B) 2730 to base sixteen C) covert CO2sixteen to base seven

OpenStudy (anonymous):

65 in base 7: \(7^2=49\) so there is a \(1\) in the seven squared place then \(65-49=16\) and there are 2 \(7\)'s in 16, to there is a \(2\) in the sevens place then \(16-14=2\) so there is a \(2\) in the ones place answer: \(122\)